OpenStudy (ujjwal):

prove that the straight lines whose direction cosines are given by the relation al+bm+cn=0 and fmn+gnl+hlm=0 are parallel if...

OpenStudy (ujjwal):

\[\sqrt{af}\pm \sqrt{bg}\pm \sqrt{ch}=0\]

OpenStudy (experimentx):

we have, al+bm+cn=0 fmn+gnl+hlm=0 l^2 + m^2 + n^2 = 1

OpenStudy (anonymous):

What is f,g,h ??

OpenStudy (ujjwal):

variables..

OpenStudy (anonymous):

They are not DC' s , right ??

OpenStudy (anonymous):

DC-->direction cosine

OpenStudy (ujjwal):

no, they are not DC's

OpenStudy (anonymous):

So. This is how I get the question. a,b,c--->DC of line 1 l,m,n-->DC of line 2 f,g,h--->Some independent variables Do correct me if I am wrong because they are essential in solving the question

OpenStudy (ujjwal):

l, m, n are direction cosines .. a,b,c are some constants and f,g,h are variables

OpenStudy (experimentx):

you sure about the if .. case??

OpenStudy (anonymous):

LOL, One set has to be DC of a line. Without that How can we prove 2 lines are parallel ??

OpenStudy (anonymous):

I meant one more set of variables

OpenStudy (ujjwal):

the question says so.. and i am sure i haven't made typing mistakes..

OpenStudy (experimentx):

\[ mn(af - gb - hc) - gc n^2 - hb m^2 = 0 \] \[ mn(2bc) + (1+b^2)m^2 + (1+c^2)n^2 - a^2 = 0 \] These two system should have a singular solution!!

OpenStudy (anonymous):

@ujjwal, It is easy. 1)Take equation al + bm + cn = 0 and write it in terms of n = ........... 2) Then substitute this n in fmn+gnl+hlm=0 3)After simplifying, divide the equation you get by m^2 Therefore you arrive at a quadratic of ax^2 + bx + c = 0 where in your case, x = l/m a,b,c-->random constants you get while solving. Now. let \[\frac{l_1}{m_1} , \frac{l_2}{m_2}\] be the two roots of the quadratic equation Now, take product of roots from quadratic equation \[\frac{l_1}{m_1} . \frac{l_2}{m_2} = -\frac{b}{a}\] where b, a are the constants you get in quadratic equation while solving Do till here , after that,I will guide you on the last part of the proving :)

OpenStudy (anonymous):

Just tell me the quadratic euqation you are getting finally

OpenStudy (anonymous):

*equation

OpenStudy (ujjwal):

In the end i get values of l1*l2, m1*m2 and n1*n2 but i need l1/l2 , m1/m2 and n1/n2.. do you get their values?

OpenStudy (anonymous):

@ujjwal , I don't require any of the values of l1*l2, m1*m2 and n1*n2 or l1/l2 , m1/m2 and n1/n2. My proof is based on different lines If you are done, can you give me the quadratic equation which is the most important part of this proof, which you get after 3rd step

OpenStudy (anonymous):

In my above comment lines means ideas

OpenStudy (ujjwal):

By a similar process i get a similar equation \[bh \left( m /n\right)^{2}+ (bg+ ch-af)\left( m/n \right)+cg=0\]

OpenStudy (anonymous):

Sorry for the delay @ujjwal . The two lines are parallel if \[\frac{l_1}{l_2} = \frac{m_1}{m_2}= \frac{n_1}{n_2}\] Therefore \frac{m_1}{n_1}= \frac{m_2}{n_2}\ Now with quadratic equation, we can say that two real and equal roots exist for this quadratic equation, therefore for quad equation of the form ax^2+bx+c=0, two roots are real and equal if \[b^2-4ac=0\] Now substitute b ,a ,c in this equation and simplify. You should get your answer

OpenStudy (anonymous):

Typo: Therefore \[\frac{m_1}{n_1}= \frac{m_2}{n_2}\]

OpenStudy (ujjwal):

thanks!

OpenStudy (anonymous):

Welcome :)